# Historically Relevant Event Structuring for Temporal Knowledge Graph Reasoning

### v1

@@ 3. Method
The overarching framework of $\mathsf{HisRES}$ is depicted in Figure 2, which follows encoder-decoder architecture and consists of three key components: (i) Multi-granularity Evolutionary Encoder3.2), which focuses on capturing correlations among concurrent facts in the most recent snapshots and modeling temporal evolution of these snapshots; (ii) Self-gating Mechanism3.3), which studies a learnable weight for each entity and adaptively merges entity representations from different encoders; (iii) Global Relevance Encoder3.4), which integrates all historically relevant facts, and aggregates their crucial neighboring semantics with a novel graph attention network.
3.1. Notations
@@
In this paper, Temporal Knowledge Graphs (TKGs) are conceptualized as sequences of subgraphs, referred to as snapshots, and denoted as $G=\{\mathcal{G}_{1},\mathcal{G}_{2},...,\mathcal{G}_{\mathcal{T}}\}$. Here, $\mathcal{T}$ represents the total number of historical timestamps, and each $\mathcal{G}_{t}$ corresponds to the collection of facts or events occurring at timestamp $t$. All historical facts correlated with query pairs can be grouped within a global graph $\mathcal{G}^{\mathcal{H}}_{t}$ in $\mathsf{HisRES}$ when making predictions, such as $(s,r,?)$ at timestamp $t$. Given $\mathcal{G}=\{\mathcal{E},\mathcal{R},\mathcal{F},\mathcal{T}\}$, where $\mathcal{E}$, $\mathcal{R}$, and $\mathcal{F}$ represent the sets of entities, relations, and facts, respectively. Each fact $f\in\mathcal{F}$ is defined as a quadruple $(s,r,o,t)$, where $s$ and $o$ belong to $\mathcal{E}$, representing the subject and object entities, $r\in\mathcal{R}$ represents the relation, and $t\in\mathcal{T}$ indicates the timestamp of the quadruple. In this paper, bold letters are used to indicate vectors. In summary, the important notations is presented in Table 1.
Table 1. The Summary of important notations.
@@
Task Definition. Our primary focus is on the extrapolation task within TKG reasoning, specifically, entity prediction in future timestamps. Given a query $q=(s,r,?,t)$ or $q=(?,r,o,t)$, where $q\in\mathcal{Q}$, our learning objective is to calculate the conditional probability of an object $o$ or subject $s$ given $(s,r,t)$ or $(r,o,t)$ and past snapshots. This is denoted as $p(o|s,r,t,\mathcal{G}_{0:t-1})$ or $p(s|r,o,t,\mathcal{G}_{0:t-1})$. It is noteworthy that we utilize object prediction as an example in our methodology section.
3.2. Multi-granularity Evolutionary Encoder
Multi-granularity evolutionary encoder serves as the cornerstone of $\mathsf{HisRES}$, which entails predicting future events using historical snapshots from the most recent past. In accordance with this principle, we consider the most neighboring $l$ snapshots as historical data and aggregate features from both structural and temporal perspectives in the multi-granularity level. It includes intra-snapshot and inter-snapshot dependencies, the latter of which can structure several adjacent snapshots simultaneously (we select every two adjacent snapshots). The illustration of one cycle in multi-granularity evolutionary encoder is depicted in Figure 3, RU denotes relation updating.
Figure 3. The achitecture of proposed multi-granularity evolutionary encoder.
3.2.1. Spatial-temporal Modeling intra-snapshot
The fundamental component of $\mathsf{HisRES}$ involves modeling spatial and temporal features from each recent snapshot, encompassing all characteristics of TKGs (i.e., entity, relation and time information), which is shown in the intra-snapshot of Figure 3.
@@ 3.2.1. Spatial-temporal Modeling intra-snapshot
Analyzing the evolution process of events, periodicity and time attributes are crucial factors. Thus, similar to previous works (Li et al., 2022b; Chen et al., 2024), we encode the time information $\Delta\textbf{t}$ based on the time interval between the current time $t_{i}$ and the predicted time $t$, as follow:
(1) $\Delta\textbf{t}=\sigma(\textbf{w}_{t}(t-t_{i})+\textbf{b}_{t}){,}$
@@
(2) $\textbf{E}^{g^{\prime}}_{t-1}=\textbf{W}_{0}([\textbf{E}^{g}_{t-1}||\Delta\textbf{t}]){,}$
@@
In the spatial aspect, the basic aggregation process resembles that of CompGCN (Vashishth et al., 2020). To merge entities and relations effectively, we employ composition operators such as ”subject+relation”, akin to RE-GCN (Li et al., 2021b). The formulation is as follows:
(3) $\textbf{o}^{(l+1)}_{t-1}=\sigma(\sum_{(s,r,o)\in\mathcal{F}_{t-1}}\textbf{W}^{(l)}_{1}(\textbf{s}^{(l)}_{t-1}+\textbf{r}^{(l)}_{t-1})+\textbf{W}^{(l)}_{2}\textbf{o}^{(l)}_{t-1})\textnormal{,}$
@@
In the temporal aspect, we incorporate the GRU cell to progressively capture evolving entity representations from each recent snapshot to fit future facts, which can be expressed as:
(4) $\textbf{E}^{g}_{t}=\mathrm{GRU}(\vec{\textbf{E}}_{t-1},\textbf{E}^{g^{\prime}}_{t-1})\textnormal{.}$
@@
Additionally, $\mathsf{HisRES}$ enriches the evolution of relations and simultaneously updates the relation representation in CompGCN during entity aggregation. The operation is as follows:
(5) $\textbf{R}^{(l+1)}_{t-1}=\sigma(\textbf{W}^{(l)}_{r}\textbf{R}^{(l)}_{t-1})\textnormal{,}$
@@
(6) $\textbf{R}_{t}=\mathrm{GRU}(\vec{\textbf{R}}_{t-1},pooling(\textbf{E}^{\mathcal{R}}_{t-1}))\textnormal{,}$
@@
3.2.2. Spatial-temporal Modeling inter-snapshot
In addition to modeling spatial-temporal correlations individually for each nearby snapshot, there are a large number of facts in TKGs that have a sequential correlation. As an example in the ICEWS dataset, (Citizen (Malaysia), Make optimistic comment, Ministry (Malaysia), t) is actually a follow-up event to (Ministry (Malaysia), Responses, Citizen (Malaysia), t-1). In this instance, the prior module designed for each individual snapshot fails to capture correlations such as this from a simple two-hop link, causing a complex mode.
To tackle this issue, we merge every two consecutive snapshots into a unified graph, which enables us to capture multi-hop association inter-snapshot using the same aggregation function as the former module (without sharing parameters). The diagram can be seen in the inter-snapshot of Figure 3. We omit the aggregation step, and denote the output of this process as $\textbf{E}^{gg^{\prime}}_{t-2:t-1}$. The evolutionary modeling proceeds as follows:
(7) $\textbf{E}^{gg}_{t}=\mathrm{GRU}(\textbf{E}^{gg^{\prime}}_{t-2:t-1},\textbf{E}^{gg}_{t-2:t-1})\textnormal{.}$
Since the graphs in this module are composed of adjacent snapshots, we do not employ relation updating as described in equation (5) of CompGCN, nor the time encoding as described in equation (2).
3.3. Self-gating Mechanism
$\mathsf{HisRes}$ incorporates a self-gating mechanism to adaptively merge multi-granularity entity embeddings from the aforementioned encoder. Given that we capture structure and evolution patterns across snapshots at various granularities, it is crucial for $\mathsf{HisRes}$ to effectively combine these features to comprehend correlations among recent snapshots. The self-gating is denoted as $\Theta\in\mathbb{R}^{|\mathcal{E}|\times d}$, which regulates a learnable weighting of different recent entity representations. Formally,
(8) $\textbf{E}_{t}=\Theta^{\mathcal{E}}\textbf{E}^{g}_{t}+(1-\Theta^{\mathcal{E}})\textbf{E}^{gg}_{t}\textnormal{,}$
(9) $\Theta^{\mathcal{E}}=\sigma(\textbf{W}_{3}\textbf{E}^{g}_{t}+\textbf{b}_{g}){,}$
@@ 3.3. Self-gating Mechanism
3.4. Global Relevance Encoder
@@
While the former modules adeptly capture the most recent evolutionary patterns from a multi-granularity perspective, it falls short in highlighting pivotal events closely related to queries within the entire historical TKGs. These events can significantly impact future trends and deserve special consideration. Unlike the widely used method of making statistics from historically related facts (Zhu et al., 2021; Li et al., 2022c; Xu et al., 2023), we incorporate an effective approach to establish a globally relevant graph that amalgamates historically relevant events. This enables explicit modeling of correlations among crucial entities and leverages comprehensive historical interactions. Furthermore, it is worth noting that our strategy, compared to recent methods (Zhang et al., 2023b; Chen et al., 2024) that construct similar global graphs to utilize historical information, achieves an even more effective capture of global correlations while requiring fewer nodes and connections.
3.4.1. Global Relevance Structuring
The idea of structuring global relevance can be seen as an expansion of historical statistics. Our method involves extracting all related facts $f_{q}\in{\mathcal{G}_{0:t-1}}$ with query pairs $(s,r)$ from the entire historical TKGs, where $q\in\mathcal{Q}_{t}$. Subsequently, the globally relevant graph $\mathcal{G}^{\mathcal{H}}_{t}$ is constructed based on these query-specific facts. This structuring process is illustrated in Figure 4.
@@ 3.4.1. Global Relevance Structuring
Figure 4. A construction illustration of the globally relevant graph. In the historical TKGs, query-specific facts are framed by green ovals, while the globally relevant graph comprises all those facts.
3.4.2. ConvGAT
@@
We consider using the fusion of multi-granularity evolutionary encoder from self-gating as the initial entity embeddings in global relevance aggregation. Since our objective of global relevance encoder is to figure out the important correlations from the whole historically relevant facts, we design a convolution-based graph attention network called ConvGAT, which integrates a 1-D convolutional operator to merge entities and relations, and an attention mechanism to focus on those crucial correlations. Firstly, we calculate the attention scores $\alpha_{o,s}$ of each edge in $\mathcal{G}^{\mathcal{H}}_{t}$, which as follows:
(10) $\theta^{(l)}_{o,s}=\frac{exp(\textbf{W}^{(l)}_{4}\sigma(\textbf{W}^{(l)}_{5}[\textbf{s}^{(l)}||\textbf{r}||\textbf{o}^{(l)}]))}{\sum_{s^{\prime}\in\mathcal{N}_{(o)}}exp(\textbf{W}^{(l)}_{4}\sigma(\textbf{W}^{(l)}_{5}[\textbf{s'}^{(l)}||\textbf{r}||\textbf{o}^{(l)}])}\textnormal{,}$
@@
where $\textbf{s},\textbf{o}\in\textbf{E}_{t}$, $\textbf{W}^{(l)}_{4}\in\mathbb{R}^{3d}$ and $\textbf{W}^{(l)}_{5}\in\mathbb{R}^{3d\times 3d}$ represent linear transformation layers, $\sigma$ corresponds to the LeakyReLU activation function, $\mathcal{N}_{(o)}$ denotes the neighboring subjects of object $o$. Notably, we do not update relation embeddings in global relevance encoder. Subsequently, the aggregation function of global relevance encoder is defined as follows:
(11) $\textbf{o}^{(l+1)}=\sigma(\sum_{(s,r,o)\in\mathcal{G}^{\mathcal{H}}_{t}}\theta^{(l)}_{o,s}\textbf{W}^{(l)}_{6}\psi(\textbf{s}^{(l)}+\textbf{r})+\textbf{W}^{(l)}_{7}\textbf{o}^{(l)})\textnormal{,}$
@@
where $\sigma$ is the RReLU activation function, $\textbf{W}^{(l)}_{6}$ and $\textbf{W}^{(l)}_{7}$ are linear transformations. The output matrix of all entity embeddings can be represented as $\textbf{E}^{\mathcal{H}}_{t}$.
Table 2. Statistics of ICEWS14s, ICEWS18, ICEWS05-15 and GDELT datasets.

### v2

@@ III Method
The overarching framework of HisRES is depicted in Figure 2, which follows encoder-decoder architecture and consists of three key modules: i) Multi-granularity Evolutionary EncoderIII-B), which focuses on capturing correlations among concurrent facts in the most recent snapshots and modeling temporal evolution of these snapshots; ii) Self-gating MechanismIII-C), which studies a learnable weight for each entity and adaptively merges entity representations from different encoders; iii) Global Relevance EncoderIII-D), which integrates all historically relevant facts, and aggregates their crucial neighboring semantics with a novel graph attention network. In DecoderIII-E), we employ a widely used ConvTransE.
Figure 2: An illustration of HisRES model architecture.
III-A Notations
@@
In this paper, Temporal Knowledge Graphs (TKGs) are conceptualized as sequences of subgraphs, referred to as snapshots, and denoted as $G=\{\mathcal{G}_{1},\mathcal{G}_{2},...,\mathcal{G}_{|\mathcal{T}|}\}$. Here, $G$ denotes the set of all TKG snapshots, $|\mathcal{T}|$ represents the total number of historical timestamps, and each snapshot $\mathcal{G}_{t}$ corresponds to the collection of facts or events occurring at timestamp $t$. Given $\mathcal{G}=\{\mathcal{E},\mathcal{R},\mathcal{F},\mathcal{T}\}$, where $\mathcal{E}$, $\mathcal{R}$, and $\mathcal{F}$ represent the sets of entities, relations, and facts, respectively. Each fact $f\in\mathcal{F}$ is defined as a quadruple $(s,r,o,t)$, where $s,o\in\mathcal{E}$, representing the subject and object entities, $r\in\mathcal{R}$ indicates the relation, and $t\in\mathcal{T}$ is the timestamp of the quadruple. Given query set $\mathcal{Q}_{t}$ at timestamp $t$, all historical facts correlated with queries like $\{q=(s,r,?,t)|q\in\mathcal{Q}_{t}\}$ before $t$, can be grouped within a global graph $\mathcal{G}^{\mathcal{H}}_{t}$ in HisRES when making predictions. Bold letters are used to indicate vectors. In summary, the important notations used in equations are presented in Table I.
TABLE I: The summary of important notations.
@@
$\omega$ hyper-parameter to control local granularity span
Task Definition. Our primary focus is on the extrapolation task within TKG reasoning, also known as link prediction problem at future timestamps.
Formal Problem Formulation. For a query $q\in\mathcal{Q}$ of the form $q=(s,r,?,t)$ or $q=(?,r,o,t)$, the formal problem formulation of HisRES is to calculate the conditional probability of an object $o$ or subject $s$ given $(s,r,t)$ or $(r,o,t)$ and historical snapshots. These probabilities are denoted $p(o|s,r,t,\mathcal{G}_{0:t-1})$ and $p(s|r,o,t,\mathcal{G}_{0:t-1})$, respectively. It is noteworthy that we utilize object prediction as the illustrative example in our methodology section (§III-F).
III-B Multi-granularity Evolutionary Encoder
Multi-granularity evolutionary encoder serves as the cornerstone of HisRES, which entails predicting future events using historical snapshots from the most recent past. In accordance with this principle, we consider the most neighboring $l$ snapshots as historical data and aggregate features from both structural and temporal perspectives in the multi-granularity span. It includes intra-snapshot and inter-snapshot dependencies, the latter of which can structure several adjacent snapshots simultaneously (we select every $\omega$ adjacent snapshots). The illustration of one cycle in multi-granularity evolutionary encoder is depicted in Figure 3.
Figure 3: The achitecture of proposed multi-granularity evolutionary encoder.
III-B1 Structural and Temporal Modeling intra-snapshot
The fundamental component of HisRES involves modeling structural and temporal features from each recent snapshot, encompassing all characteristics of TKGs (i.e., entity, relation and time information), which is shown in the intra-snapshot of Figure 3.
@@ III-B1 Structural and Temporal Modeling intra-snapshot
Analyzing the evolution process of events, periodicity and time attributes are crucial factors. Thus, similar to previous efforts [42, 16], we encode the time information $\Delta\textbf{t}$ based on the time interval between the current time $t_{i}$ and the predicted time $t$, as follow:
$\Delta\textbf{t}=\sigma(\textbf{w}_{t}(t-t_{i})+\textbf{b}_{t}){,}$ (1)
@@
$\textbf{E}^{g^{\prime}}_{t-1}=\textbf{W}_{0}([\textbf{E}^{g}_{t-1}||\Delta\textbf{t}]){,}$ (2)
@@
In the structural aspect, the basic aggregation resembles that of CompGCN [28]. To merge entities and relations effectively, we employ composition operators such as ”subject+relation”, akin to RE-GCN [6]. The formulation is as follows:
$\textbf{o}^{(l+1)}_{t-1}=\sigma(\sum_{(s,r,o)\in\mathcal{F}_{t-1}}\textbf{W}^{(l)}_{1}(\textbf{s}^{(l)}_{t-1}+\textbf{r}^{(l)}_{t-1})+\textbf{W}^{(l)}_{2}\textbf{o}^{(l)}_{t-1})\textnormal{,}$ (3)
@@
In the temporal aspect, we incorporate a Gated Recurrent Unit (GRU) cell to progressively capture evolving entity representations from each recent snapshot to fit future facts, which can be expressed as:
$\textbf{E}^{g}_{t}=\mathrm{GRU}(\vec{\textbf{E}}_{t-1},\textbf{E}^{g^{\prime}}_{t-1})\textnormal{.}$ (4)
@@
Additionally, HisRES enriches the evolution of relations and simultaneously updates the relation representation in CompGCN during entity aggregation. The operation is as follows:
$\textbf{R}^{(l+1)}_{t-1}=\sigma(\textbf{W}^{(l)}_{r}\textbf{R}^{(l)}_{t-1})\textnormal{,}$ (5)
@@
$\textbf{R}_{t}=\mathrm{GRU}(\vec{\textbf{R}}_{t-1},pooling(\textbf{E}^{\mathcal{R}}_{t-1}))\textnormal{,}$ (6)
@@
III-B2 Structural and Temporal Modeling inter-snapshot
In addition to modeling structural and temporal correlations individually for each nearby snapshot, there are a large number of facts in TKGs that have a sequential correlation. As an example in the ICEWS dataset, (Citizen (Malaysia), Make optimistic comment, Ministry (Malaysia), t) is actually a follow-up event to (Ministry (Malaysia), Responses, Citizen (Malaysia), t-1). In this instance, the prior module designed for each individual snapshot fails to capture correlations such as this from a simple two-hop link, causing a complex mode.
To tackle this issue, we merge every $\omega$ consecutive snapshots into a unified graph, which enables us to capture multi-hop association inter-snapshot using the same aggregation function as the former module (without sharing parameters). The diagram can be seen in the inter-snapshot of Figure 3. We omit the aggregation step, and denote the output of this process as $\textbf{E}^{gg^{\prime}}_{t-2:t-1}$. The evolutionary modeling proceeds as follows:
$\textbf{E}^{gg}_{t}=\mathrm{GRU}(\textbf{E}^{gg^{\prime}}_{t-2:t-1},\textbf{E}^{gg}_{t-2:t-1})\textnormal{.}$ (7)
Notably, inter-snapshot modeling involves a sequence of composite graphs spanning larger time intervals with multiple timestamps. Consequently, the time encoding used to represent the time gap relative to the predicted time is not suitable for this module.
III-C Self-gating Mechanism
HisRES incorporates a self-gating mechanism to adaptively merge multi-granular embeddings from the aforementioned encoder. Given that we capture structure and evolution patterns across snapshots at various granularities, it is crucial for HisRES to effectively combine these features to comprehend correlations among recent snapshots. The self-gating is denoted as $\Theta\in\mathbb{R}^{|\mathcal{E}|\times d}$, which regulates a learnable weighting of different recent entity representations. Formally,
$\textbf{E}_{t}=\Theta^{\mathcal{E}}\textbf{E}^{g}_{t}+(1-\Theta^{\mathcal{E}})\textbf{E}^{gg}_{t}\textnormal{,}$ (8)
$\Theta^{\mathcal{E}}=\sigma(\textbf{W}_{3}\textbf{E}^{g}_{t}+\textbf{b}_{g}){,}$ (9)
@@ III-C Self-gating Mechanism
III-D Global Relevance Encoder
@@
While the former modules adeptly capture the most recent evolutionary patterns from a multi-granularity perspective, it falls short in highlighting pivotal events closely related to queries within the entire historical TKGs. These events can significantly impact future trends and deserve special consideration. Unlike the widely used method of making statistics from historically related facts [8, 7, 9], we propose an effective approach to establish a globally relevant graph that amalgamates historically relevant events. This enables explicit modeling of correlations among crucial entities and leverages comprehensive historical interactions. Furthermore, it is worth noting that our strategy, compared to recent methods [15, 16] that construct similar global graphs to utilize historical information, achieves an even more effective capture of global correlations while requiring fewer nodes and connections.
@@
Figure 4: An illustrative construction of the globally relevant graph. In the historical TKGs, query-specific facts are framed by green ovals, while the globally relevant graph is structured by all those facts.
III-D1 Global Relevance Structuring
The idea of structuring global relevance can be seen as an expansion of historical statistics. It involves extracting all related facts $f_{q}\in{\mathcal{G}_{0:t-1}}$ with query pairs $(s,r)$ from the entire historical TKGs, where $q\in\mathcal{Q}_{t}$. Subsequently, the globally relevant graph $\mathcal{G}^{\mathcal{H}}_{t}$ is constructed based on these query-specific facts. This structuring process is illustrated in Figure 4. For queries $(e_{1},r_{1},?,t)$ and $(e_{3},r_{3},?,t)$, HisRES first retrieves relevant facts from all historical snapshots (denoted as ${\mathcal{G}_{0}:\mathcal{G}_{t-1}}$), such as ${(e_{1},r_{1},e_{4})\in\mathcal{G}_{0}}$, ${(e_{1},r_{1},e_{2})\in\mathcal{G}_{1}}$, and ${(e_{3},r_{3},e_{1})\in\mathcal{G}_{t-2}}$. These facts are then integrated into the globally relevant graph.
III-D2 ConvGAT
@@
We consider using the fusion of multi-granularity evolutionary encoder from self-gating as the initial entity embeddings in global relevance aggregation. Since our objective of global relevance encoder is to figure out the important correlations from the whole historically relevant facts, we design a convolution-based graph attention network called ConvGAT, which integrates a 1-D convolutional operator to merge entities and relations, and an attention mechanism to focus on those crucial correlations.
Firstly, for each link in $\mathcal{G}^{\mathcal{H}}_{t}$, we calculate the attention scores $\alpha_{o,s}$ based on the embeddings of quadruples, which as follows:
$\theta^{(l)}_{o,s}=\frac{exp(\textbf{W}^{(l)}_{4}\sigma(\textbf{W}^{(l)}_{5}[\textbf{s}^{(l)}||\textbf{r}||\textbf{o}^{(l)}]))}{\sum_{s^{\prime}\in\mathcal{N}_{(o)}}exp(\textbf{W}^{(l)}_{4}\sigma(\textbf{W}^{(l)}_{5}[\textbf{s}^{\prime(l)}||\textbf{r}||\textbf{o}^{(l)}])}\textnormal{,}$ (10)
@@
where $\textbf{s},\textbf{o}\in\textbf{E}_{t}$, $\textbf{r}\in\textbf{R}_{t}$, $\textbf{W}^{(l)}_{4}\in\mathbb{R}^{3d}$ and $\textbf{W}^{(l)}_{5}\in\mathbb{R}^{3d\times 3d}$ represent linear transformation layers, $\sigma$ corresponds to the LeakyReLU activation function, $\mathcal{N}_{(o)}$ denotes the neighboring subject set of object $o$. Notably, we do not update relation embeddings in global relevance encoder. Subsequently, for each link in $\mathcal{G}^{\mathcal{H}}_{t}$, the aggregation function of global relevance encoder is defined as follows:
$\textbf{o}^{(l+1)}=\sigma(\sum_{(s,r,o)\in\mathcal{G}^{\mathcal{H}}_{t}}\theta^{(l)}_{o,s}\textbf{W}^{(l)}_{6}\psi(\textbf{s}^{(l)}+\textbf{r})+\textbf{W}^{(l)}_{7}\textbf{o}^{(l)})\textnormal{,}$ (11)
@@
where $\sigma$ is the RReLU activation function, $\textbf{W}^{(l)}_{6}$ and $\textbf{W}^{(l)}_{7}$ are linear transformations, $\psi$ denotes a 1-d convolution operator for the fusion of entity and relation embeddings. The output matrix of all entity embeddings from global relevance encoder can be represented as $\textbf{E}^{\mathcal{H}}_{t}$.
III-E Decoder
To achieve comprehensive predictions by leveraging entity and relation information from TKGs, HisRES adopts ConvTransE [25] as the score function, a widely used decoder in TKG reasoning, defined as follows:
$p(o|s,r,t,\mathcal{G}_{0:t-1})=softmax(\mathrm{ConvTransE}(\textbf{s}^{\phi},\textbf{r}){\textbf{E}^{\phi}_{t}})\textnormal{,}$ (12)
$\textbf{E}^{\phi}_{t}=\Theta^{\mathcal{H}}\textbf{E}^{\mathcal{H}}_{t}+(1-\Theta^{\mathcal{H}})\textbf{E}_{t}{,}$ (13)
$\Theta^{\mathcal{H}}=\sigma(\textbf{W}_{8}\textbf{E}^{\mathcal{H}}_{t}+\textbf{b}_{gg}){,}$ (14)
III-F Learning Objective
Entity prediction in KGs can be regarded as a multi-classification task. Hence, HisRES employs Cross-Entropy as the loss function. Importantly, we optimize the relation prediction to enhance relation representations, akin to TiRGN [7] and LogCL [16], to further enhance the performance of entity prediction. In addition, a coefficient $\alpha$ is introduced to control the relative weight of the two tasks. The learning objective is formalized as follows:
$\mathcal{L}$ $=\alpha\sum_{(s,r,t)\in\mathcal{Q}^{e}_{t}}y^{e}_{t}\log p(o|s,r,t,\mathcal{G}_{0:t-1})$ (15)
III-G Complexity Analysis
We conduct complexity analysis for the proposed HisRES to showcase the efficiency and our trade-offs in capturing multi-granular and global relevant interactions. Note that the embedding dimension $d$ is omitted from the notation. The time complexity of the multi-granularity evolutionary encoder is $O(|2l-\omega+1|(|\mathcal{E}|+|\mathcal{R}|))$, as we consider both intra- and inter-snapshot correlations simultaneously. The time complexity of the global relevance encoder is $O(|\mathcal{E}|+2|\mathcal{F}_{\mathcal{Q}t}|)$, where $|\mathcal{F}_{\mathcal{Q}_{t}}|$ denotes the number of query-specific facts in $\mathcal{G}^{\mathcal{H}}_{t}$. Therefore, the total computational complexity can be simplified to $O(|2l-\omega|(|\mathcal{E}|+|\mathcal{R}|)+2|\mathcal{F}_{\mathcal{Q}_{t}}|)$. This analysis demonstrates that HisRES achieves outstanding performance without a significant increase in computational complexity. The computational complexity of HisRES is compared with baseline methods in Table II. Here, $l$ represents the snapshot length, $\mathcal{M}_{c}$ denotes the convolution operations in TiRGN, while $P$ and $4mC$ correspond to the attention mechanism and contrastive learning components in LogCL, respectively. The additional computational overhead primarily originates from two aspects: the multi-granularity processing (i.e., $|l-\omega|$) and the global relevance structuring (i.e., $|\mathcal{F}_{\mathcal{Q}_{t}}|$). Although this approach involves comprehensive historical information and intensive correlation analysis, it remains computationally tractable and effectively serves the task objectives.
TABLE II: Computational Complexity Comparison of Baselines with HisRES.
Model Computational Complexity
TiRGN $\mathcal{O}(|l|(|\mathcal{E}|+|\mathcal{R}|+\mathcal{M}_{c})+|\mathcal{T}|)$
LogCL $O(|l|(|\mathcal{E}|+|\mathcal{R}|+P)+2|\mathcal{F}||\mathcal{T}|+4mC)$
HisRES $O(|2l-\omega|(|\mathcal{E}|+|\mathcal{R}|)+2|\mathcal{F}_{\mathcal{Q}_{t}}|)$
IV Experiments
TABLE III: Statistics of ICEWS14s, ICEWS18, ICEWS05-15 and GDELT datasets.